### v1

# Coexistence of Anderson Localization and Quantum Scarring in Two Dimensions

Abstract
We study finite two-dimensional disordered systems with periodic confinement. At low energies, eigenstates exhibit strong Anderson localization, while at higher energies a subset of states forms variational scars with anisotropic intensity patterns that violate random wave expectations. Scaling theory predicts that all states localize in two dimensions, yet energy-dependent localization lengths and finite system size allow these regimes to coexist. We demonstrate that this coexistence produces distinct, robust signatures in both spatial intensity patterns and spectral statistics that are directly observable in mesoscopic electronic, photonic, and cold atom systems.
Introduction– In disordered quantum systems, the breakdown of ergodicity can manifest in several distinct forms, from Anderson localization (AL) in non-interacting settings to weakly non-ergodic structures that defy random‑wave expectations [DeTomasi2020Multifractality, Altshuler2016Nonergodic, Das2025Emergent, Xu2023NEE]. Anderson’s original insight is that quenched disorder can localize single-particle eigenstates and suppress transport. This principle establish an essential tenet of mesoscopic physics [Anderson1958, Abrahams1979, Billy2008, Roati2008]. In parallel, deviant wavefunctions that concentrate along classical periodic orbits (POs), known as quantum scars [Heller_book], offer a different route to non-ergodic behavior within single-particle systems. Such scarred states are reported in billiard-like platforms [Heller_phys.rev.lett_53_1515_1984, Noeckel1997, Bogomolny2012] as well as in quantum wells [Luukko_sci.rep_6_37656_2016, keski-rahkonen_phys.rev.lett_123_214101_2019, Selinummi2024, Chalangari2025, keskirahkonen_phys.rev.e_112_L012201_2025], and recently, scarred eigenstates were directly visualized in a graphene-based quantum well [Ge_nature_635_841_2024]. Many-body analogs of these scars have also been uncovered in constrained lattice models [Turner2018, Serbyn2021]. These phenomena are nowadays routinely probed across platforms with high control, such as two‑dimensional (2D) electron gases, disordered photonic media, waveguide arrays, and ultracold atoms  [Schwartz2007, Lahini2008, Kondov2011, directVis_Ge2024, PhysRevB103104201, White2020].
In 2D disordered systems, the localization length $\xi(E)$ grows rapidly with energy $E$, such that finite samples with linear size $L$ may realize both $L\ll\xi(E)$ and $L\gg\xi(E)$ regimes [RMTinQT]. In such systems, high-energy states may appear delocalized and exhibit Wigner–Dyson statistics [Mehta_book, PhysRevB103104201], commonly regarded as a quantum hallmark of chaos [Stockmann_book]. Yet scaling theory [Abrahams1979] excludes a true metallic phase, as in non-interacting electronic systems with time-reversal symmetry, all states localize in the thermodynamic limit for any strength of uncorrelated disorder. Short-range correlations in the disorder do not, in general, invalidate the scaling argument, but instead renormalize the effective disorder strength and can modify $\xi(E)$ by orders of magnitude [Conley2014, Monsarrat2022]. Nevertheless, many aspects of AL are still poorly understood [Anderson_localization_book], particularly in the continuum limit and in the presence of long-range correlations.
In bulk-like continuous systems, low-energy states exhibit strong backscattering and AL, while at higher energies smooth, non-integrable disorder generates semiclassical caustics associated with branched flow [Daza2021, Heller2021, Heller_book], observed in, e.g., 2D electron gases [Topinka2001] and optical media [laserbranch].
Moreover, in finite systems whose bounded geometries support families of classical POs, these orbits can play an enhanced role due to the repeated returns of semiclassical trajectories. Local disorder lifts the degeneracies of the underlying periodic potential, generating small clusters of near-degenerate eigenstates associated with classical POs. Coherent interference within these clusters produces intensity enhancement along the corresponding POs, giving rise to variational scars [Luukko_sci.rep_6_37656_2016]. The orientation of these scarred states tends to be selected variationally, either maximizing or minimizing the overlap with the effective impurity landscape [simo, keski-rahkonen_phys.rev.b_97_094204_2017, keski-rahkonen_phys.rev.lett_123_214101_2019, Luukko_phys.rev.lett_119_203001_2017]. These features signal a form of weak ergodicity breaking that is distinct from typical localization phenomena, such as Anderson, dynamical or many-body localization.
Despite the progress on these frontiers, the manner in which finite-size effects and continuum disorder conspire to yield coexisting non-ergodic regimes across the spectrum in realistic systems remains elusive. To address this issue, we here investigate a finite 2D continuum Hamiltonian with prototypical disorder of Gaussian potential bumps serving as a proxy for local impurities. Our analysis reveals regimes that separate across the energy spectrum: (i) conventional AL eigenstates dominating the low-energy sector, (ii) genuinely ergodic delocalized states following random-matrix-theory statistics in the sense of the Bohigas-Giannoni-Schmit conjecture [Bohigas_phys.rev.lett_52_1_1984], while (iii) a set of high-energy scarred states exhibiting intensity patterns that deviate from random-wave predictions of the Berry conjecture [Berry_j.phys.a_10_2083_1977]. Representative eigenstates of these classes are shown in Fig. 1 (upper panel and in Supplemental Material (SM) S1), illustrating a spectrum of behaviors ranging from AL [Fig. 1(a)], to ergodic delocalization [Fig. 1(b)] and varying degrees of scarring [Figs. 1(c & d)]. These regimes are distinguished through an analysis based on complementary observables, including spatial profiles, inverse participation ratios (IPR), multifractal spectra, and deviations from universal level statistics.
Model Setting– We consider a 2D potential profile $V_{\mathrm{ext}}(\mathbf{r})$, confined by Dirichlet hard-wall boundaries to a region of size $L\times L$, where $L$ denotes the number of quantum-wells along each direction, together with Gaussian impurities $V_{\text{imp}}(\mathbf{r})$. By utilizing the imaginary-time propagation method [LuukkoP.J.J.2013Itpc], we obtain the lowest $\sim 4\times 10^{3}-10^{4}$ eigenstates of the time-independent Schrödinger equation $H\psi_{n}(\mathbf{r})=E_{n}\psi_{n}(\mathbf{r})$ with
$H=-\frac{1}{2}\nabla^{2}+V_{\mathrm{ext}}(\mathbf{r})+V_{\mathrm{imp}}(\mathbf{r}),$ (1)
in atomic units (a.u.). In the following, eigenenergies $E_{n}$ are normalized as $\tilde{E}_{n}=(E_{n}-E_{\min})/(E_{\max}-E_{\min})$, where $E_{\min}$ is the energy of the ground state and $E_{\max}$ is the highest eigenenergy obtained in the numerics.
We focus on non-integrable Fermi wells that form the backbone of external potential $V_{\text{ext}}$. The corresponding classical dynamics are chaotic [Daza2021, klages_phys.rev.lett_122_064102_2019, graf_entropy_26_492_2024, Toivonen2025], ensuring that no special symmetries or separable structure constrain the dynamics. The system consists of a periodic array of circular wells described by the symmetrized Fermi profile
$V_{\text{ext}}(\mathbf{r})=V_{0}\frac{\coth{\Big(\frac{r_{0}}{2d}\Big)}\sinh(r_{0}/d)}{\cosh(r/d)+\cosh(r_{0}/d)},$ (2)
where $r_{0}$, $d$, and $V_{0}$ set the well radius, edge steepness, and well-depth [stefanPRB], respectively. Throughout this work, we fix the Fermi potential parameters to $r_{0}=0.8$, $d=0.03$, $V_{0}=20$, and spatial period $a=2$ (all in a.u.), chosen to qualitatively match earlier experimental conditions [fermiexp1, fermiexp2].
Figure 1: Energy vs Disorder. (upper panel) Representative eigenstates across disorder-energy parameter space: (a) Anderson-localized, (b) delocalized, (c) scarred, (d) weakly scarred. (lower panel) Disorder strength $\langle A\rangle/V_{0}$ plotted against independently normalized energy $\tilde{E}_{n}$], with the colormap indicating $\log_{10}(\mathrm{IPR}_{2})$. The gray vertical dashed-line marks the scaled well-depth at 0.08 on the normalized energy axis. Simulations use fixed parameters $r_{0}=0.8$, $d=0.03$, $V_{0}=20$, $a=2$, and $L=5$ (all in a.u.).
Figure 2: Mesoscopic scaling analysis. (a) Normalized energy $\tilde{E}_{n}$ versus system size L=413L=4-13 with identical impurity density 0.4, and fixed disorder strength $\langle A\rangle/V_{0}=0.3$. The color scale shows $\log_{10}(\mathrm{IPR_{2}})$. (b) Log–log plot of $\mathrm{IPR}_{2}$ versus $L$ for the same systems, with color indicating the normalized eigenenergy. Overlaid symbols show the mean scaling of Anderson-localized (green), delocalized (red), and scarred (yellow) states.
Disorder is introduced via randomly distributed Gaussian bumps as
$V_{\text{imp}}(\mathbf{r})=\sum_{i}A_{i}e^{-|\mathbf{r}-\mathbf{r}_{i}|^{2}/(2\sigma_{i}^{2})},$ (3)
with amplitude $A_{i}$, width $\sigma_{i}$, and position $\mathbf{r}_{i}$, which is a standard approach for modeling imperfections or impurities in quantum wells [hirose_phys.Rev.B_63_075301_2001, hirose_phys.rev.B_65_193305_2002]. The bump positions $\mathbf{r}_{i}$ are sampled independently from a uniform distribution, yielding an average density $\rho$ per unit area. The average impurity width is fixed in all simulations at $\langle\sigma_{i}\rangle=0.2$. The dimensionless ratio $\langle A\rangle/V_{0}$, where $\langle A\rangle$ is the average bump amplitude, characterizes the overall disorder strength.
Coexisting non-ergodic regimes– We begin by assessing the spectral and wavefunction properties of the system through the generalized IPR for a normalized eigenfunction $\psi(\mathbf{r})$ as
$\mathrm{IPR}_{q}=\int d^{2}\mathbf{r}\,|\psi(\mathbf{r})|^{2q}\sim L^{-D_{q}(q-1)}\,\quad(q\geq 2),$ (4)
which subsequently defines generalized fractal dimension $D_{q}$ [Girvin_Yang_2019]. We here restrict attention to the standard second-order moment of $q=2$, which is used as a quantitative benchmark to classify the different spectral regimes.
Figure 1 (lower panel) illustrates a typical localization–delocalization crossover in the disorder–energy plane for a system with fixed size $L$ and impurity density $\rho$. The colormap shows $\log_{10}(\mathrm{IPR}_{2})$, and the gray vertical dashed-line marks the scaled well-depth energy $V_{0}$. At energies below and near the well-depth, $E_{n}\lesssim V_{0}$, eigenstates are dominated by potential confinement (see a representative example in Fig.1(a)), and increasing disorder strength $\langle A\rangle/V_{0}$ disrupts inter-well resonances. This results in a broadening of the energy window populated by strongly AL states. At energies $E_{n}\gtrsim V_{0}$, the IPR landscape exhibits two quantitatively distinct regimes. One consists of very low IPR eigenstates in which scarring is essentially absent [see Fig. 1(b)]. For these states the localization length exceeds the system size, and de Broglie wavelength $\lambda(E)\sim 2\pi/\sqrt{2(E-\langle V\rangle)}$ $\langle V\rangle$ denotes the expectation value of total potential energy] is still not small enough to resolve the underlying periodic structure.
Upon further increasing the energy, de Broglie wavelength becomes sufficiently short to resolve the periodic structure of the confinement. In this regime, scarred states emerge as anisotropic outliers embedded within the delocalized background, corresponding to variational scarring. As the disorder strength increases, both the low- and intermediate- energy windows [i.e., Fig. 1(a) and (b)] broaden, pushing the scar-bearing region to progressively higher energies. In strongly disordered cases, the remaining scars are weakened and occupy a diminishing fraction of the accessible spectrum, consistent with previous observations [KeskiRahkonen2019b, simo].
Figure 2(a) shows how the spatial structure of the eigenstates evolves with respect to system size for fixed disorder strength $\langle A\rangle/V_{0}$ and impurity density $\rho$. For the scaling analysis, we compute a fixed set of $10^{4}$ eigenstates for each system size. At energies $E_{n}\lesssim V_{0}$, AL states occupy an increasingly large fraction of the sampled spectrum as $L$ increases. This trend is consistent with the scaling theory, which predicts that at fixed energy increasing $L$ drives more states into the regime $L/\xi(E)\gg 1$.
As discussed in Fig. 1, at energies $E_{n}\gtrsim V_{0}$, the spectrum contains both delocalized and scarred eigenstates. In 2D, the mean spacing between adjacent energy levels scales as $\sim L^{-2}$ [Girvin_Yang_2019]. Consequently, a fixed set of $10^{4}$ computed eigenstates spans a progressively narrower absolute energy window as the system size increases (see Fig. S3 in SM). Therefore, larger systems require increasingly greater number of eigenstates to reach energies at which the de Broglie wavelength becomes short enough to resolve the underlying geometry. The fixed sampled spectrum is thus dominated by delocalized states, while the scar-bearing window are pushed to higher relative energies.
The coexistence of AL, delocalized, and scarred states within the same sample can thus be explained by the interplay between the energy-dependent $\xi(E)$ and the finite system size. We find that scarred states emerge when (i) $\xi(E)$ becomes comparable to or exceeds $L$, and (ii) de Broglie wavelength becomes sufficiently small to allow eigenfunctions to resolve the underlying PO structure imposed by the bounded geometry. In this regime, the system admits families of unstable and marginally stable POs embedded within the chaotic sea, and interference along these trajectories imprints the characteristic directional, scarred patterns onto the corresponding eigenstates [kaplan_ann.phys_264_171_1998, Heller_book].
Figure 2(b) exhibits the same set of eigenstates as in Fig. 2(a), now presented in a log–log plot of the size $L$ versus the localization measure $\mathrm{IPR}_{2}$. The colormap encodes the normalized eigenenergy, enabling the different localization regimes to be visually distinguished. Several classes of states emerge: some exhibit integer effective fractal dimensions, corresponding to scaling laws $\mathrm{IPR}_{2}\propto\xi^{-2},\;L^{-2}$, or $L^{-1}$, while others display non-integer multifractal scaling. The mean scaling of AL states (green markers) and delocalized ergodic states (red markers) agrees reasonably well with the expected universal power-laws, with moderate deviations in small systems attributable to finite-size effects [Girvin_Yang_2019]. Scarred states (yellow markers) behave differently. For small systems, their mean scaling appears approximately linear on the log–log plot, however, this trend breaks down as $L$ increases. For a fixed number of the computed eigenstates, the averaged scaling of scarred states drifts toward the delocalized branch, reflecting that variational scars in this energy range rely on finite-size periodic structure. As the system grows, these orbits carry diminishing spectral weight and the corresponding scars become increasingly masked by the growing density of competing states. This does not preclude the existence of strong high-energy scars (such as superwire-like modes [Daza2021, graf_entropy_26_492_2024]) in larger periodic systems, rather, it indicates that these modes shift to higher energies outside the numerically sampled window.
Figure 3: Level-spacing statistics. (a) Distribution of consecutive level-spacing ratios for disordered systems of sizes $L=4$ and $L=13$. The corresponding symmetrized gap-ratio averages are $\langle\tilde{s}\rangle\approx 0.389$ and $0.524$, respectively. (b) Comparison of weak and strong disorder strengths, $\langle A\rangle/V_{0}=0.1$ and $2$, for $L=5$, yielding $\langle\tilde{s}\rangle\approx 0.388$ and $0.525$, respectively.
Figure 4: Quantum vs Classical. (a) Ratio of the kinetic and potential energy expectation values $\langle T\rangle/\langle V\rangle$ as a function of normalized energy in the clean system. High-energy eigenstates show anomalously large ratios, indicating the kinetic-dominated precursors of scarred modes. (b) Upon introducing disorder, degeneracy lifting generates quasi–one-dimensional scarred states that continue to exhibit elevated $\langle T\rangle/\langle V\rangle$ compared to delocalized, fractal, or ergodic states. Insets illustrate real-space probability densities of the selected eigenstates.
For localized eigenstates, the angularly averaged density decays exponentially, $|\psi(r)|^{2}\propto e^{-2r/\xi_{\mathrm{tail}}}$ (see Fig. S2 in SM), consistent with conventional Anderson localization. Scarring also leaves clear signatures in level-spacing statistics. We quantify spectral correlations employing the ratio of consecutive energy spacings, i.e., $s_{n}=\delta_{n}/\delta_{n}-1$, where $\delta_{n}=(E_{n+1}-E_{n})$. This statistic does not require unfolding [Atas2013] and cleanly resolves the Poisson behavior $P_{\mathrm{Poisson}}(s)\propto 1-2s$ from the random matrix theory result $P_{\mathrm{GOE}}(s)\propto s$ given by the Gaussian orthogonal ensemble (GOE) which is the Wigner–Dyson distribution for time-reversal–symmetric systems.  [^1] For completeness, the exact distributions for the Poisson and GOE limits are $P_{\mathrm{Poisson}}(s)=1/(1+s)^{2}$ and $P_{\mathrm{GOE}}(s)=\dfrac{27}{8}\,\dfrac{s+s^{2}}{(1+s+s^{2})^{5/2}}$, respectively. At large values of $s$ ($s\gg 1$), both distributions scales as $P(s)\propto s^{-2}$. However, the key distinction lies in level repulsion, i.e., the tendency of adjacent energy levels to avoid clustering, which manifests at small values of $s$. In addition to $P(s)$, we characterize spectral correlations using the symmetrized ratio $\tilde{s}_{n}=\min\left(s_{n}\;,1/s_{n}\right)\in(0,1)$, whose mean value provides a compact, unfolding-free diagnostic of spectral statistics [Atas2013]. In Poisson limit $\langle\tilde{s}\rangle_{\mathrm{Poisson}}\approx 0.386$, whereas for GOE $\langle\tilde{s}\rangle_{\mathrm{GOE}}\approx 0.536$.
Figs. 3(a) and (b) summarize the evolution of the level-spacing statistics under disorder. One expects that strong disorder as well as large system size drive the level statistics toward the Poisson distribution. In contrast, our results uncover the opposite trend, in agreement with the IPR analysis. As shown in Fig. 3(a), for small systems ($L=4$), the distribution is close to Poissonian. Such behavior is consistent with weak level correlations caused by localized eigenstates and spectral clustering, which suppress level repulsion. Similar effects in the presence of perturbation-induced scarring have been reported previously in single 2D quantum wells [KeskiRahkonen2019b].
As the system size increases ($L=13$), the statistics shift toward the Wigner-Dyson form. This indicates that the contribution of scarred states within the relevant spectral window is reduced, even though the total number of AL states continues to grow. Increasing the disorder strength [Fig. 3(b)] produces a similar crossover, where the disappearance of scar-related signatures become apparent in the level statistics. Therefore, for a relevant choice of spectral window and number of levels, level-spacing statistics can provide indications of the presence of scarred states.
These results differ from standard tight-binding predictions and thus warrant clarification. In the deep-well limit, where each well hosts a single impurity and Gaussian potentials do not overlap, the disorder reduces to uncorrelated on-site randomness and the continuum model approaches the Anderson tight-binding description (see SM for pedagogical derivation). Only in this regime is a direct comparison meaningful. In finite tight-binding systems, the localization length $\xi$ may exceed the system size $L$, leading to apparently delocalized states that localize asymptotically. By contrast, smooth confinement and correlated disorder in the continuum support interference effects and scar-like high-energy morphologies that are strongly suppressed in coarse-grained lattice descriptions.
To clarify the origin of these continuum-only features, we compare the quantum eigenstates with their classical skeleton. At energies above the confinement barrier, classical trajectories produce smoothly varying fluctuations in ratio of kinetic $T$ to potential $V$ energy. However, Fig. 4(a) unveils that quantum eigenstates of the clean system deviate from this classical expectation even at energies far above the well-depth. They exhibit pronounced fluctuations in $\langle T\rangle/\langle V\rangle$, forming extended stripe-like patterns, reflecting the underlying potential periodicity. These kinetic energy dominated states act as precursors of scars.
When disorder is introduced [Fig. 4(b)], the near-degenerate clusters of high-energy Bloch-like states are lifted, causing their coherent interference to collapse the stripe precursors into sharply localized, quasi–one-dimensional channels, displayed in the inset of Fig. 4(b) [see also Figs. 1(c & d)]. These kinetic-dominated outliers correspond to scarred modes, arising from interference within groups of nearly degenerate states. The resulting channels are selected variationally, as they extremize the overlap of the eigenstates with the effective impurity landscape, which in a periodic potential is dominated by impurities inside the wells where the wavefunction has appreciable weight. Their persistence under disorder underscores their quantum-interference origin. Rather than fully delocalizing, the high-energy states retain a memory of the potential periodicity and form anisotropic structures absent in coarse-grained Anderson models. This mechanism accounts for the joint effect of smooth correlated disorder and periodic confinement in producing non-ergodic morphologies at energies for which classical intuition anticipates featureless, ergodic dynamics.
Summary and Discussion– We have demonstrated the coexistence of Anderson-localized, delocalized, and variationally scarred states in finite 2D disordered continua, arising from finite-size crossovers governed by the rapid growth of the localization length with energy. Although scaling theory rules out a true metallic phase in the thermodynamic limit for short range hoppings, the combination of short but smooth correlated disorder and an underlying periodic confinement produces quasi-localized, anisotropic high-energy modes that have no analogue in typical coarse-grained tight-binding models owing to truncation.
Rather than a sharp transition, the apparent splitting of spectral ergodicity reflects the interplay between localization length and the system size, which allows all three types of states to coexist within the same spectral window. The high-energy stripe-like structures present in the clean system collapse into quasi–one-dimensional channels once correlated disorder is introduced. Interference among disorder-induced near-degenerate states leads to directional variational scars unique to finite, periodically structured continua. Collectively, these results demonstrate the coexistence of conventional Anderson localization and weak ergodicity breaking in realistic mesoscopic systems, with experimentally accessible signatures in electronic, photonic, and cold-atom platforms.
Acknowledgments.– We acknowledge CSC — Finnish IT Center for Science for computational resources, and Research Council of Finland, ManyBody2D Project No. 349956, for financial support. This project is also supported by the National Science Foundation (Grant No. 2403491).
Coexistence of Anderson Localization and Quantum Scarring in Two Dimensions
Fartash Chalangari , Anant Vijay Varma , Joonas Keski-Rahkonen and Esa Räsänen
(Supplementary Material)
In this Supplementary we provide examples of different ergodic and non-ergodic states. Details of the Anderson state tail analysis, comparison of clean and disorder periodic systems with different sizes in a typical scenario. Moreover, for pedagogy a tight-binding derivation assuming one narrow Gaussian bump in each potential well.
Figure S1: Typical Spectrum hosting multiple species of eigenstates. Selected eigenstates of a system with $L=5$. White circles mark periodic wells and surrounding gray regions indicate confinement barriers. (a) and (c) are respectively the typical low and high energy eigenstates when $V_{\mathrm{imp}}(x,y)=0$ (b) & (d) are eigenstates with same eigen indices after inclusion of Gaussian bumps ($V_{\mathrm{imp}}(x,y)$ is non zero).
=======\!=\!=\!=\!=\!=\; [I] Tail Analysis
For an Anderson localized state, the density can be written as Eq. S-1, and to extract the localization length, we fit $\ln|\psi(r)|^{2}$ versus r only in the linear tail window where $d[\ln A(r)]/dr\approx 0$. In this window, the prefactor is effectively constant, $A(r)=C$, and from normalization condition of eigenstates, we can derive the relation between $\text{IPR}_{2}$ and localization length:
$\begin{split}&1=\int|\psi(r)|^{2}\;d^{2}r=\int_{0}^{\infty}2\pi C\;r\;e^{-r/\xi}\;dr\\ &\Rightarrow 2\pi C\int_{0}^{\infty}re^{-r/\xi}\;dr=2\pi C\;\xi^{2}=1\\ &\Rightarrow C=\frac{1}{2\pi\xi^{2}}\end{split}$ (S-1)
$\begin{split}&\text{IPR}_{2}=\int|\psi(r)|^{4}\;d^{2}r=\int_{0}^{\infty}2\pi r\;C^{2}e^{-2r/\xi}\;dr\\ &\Rightarrow\text{IPR}_{2}=2\pi\left(\frac{1}{4\pi^{2}\xi^{4}}\right)\left(\frac{\xi^{2}}{4}\right)=\frac{1}{8\pi}\xi^{-2}\end{split}$ (S-2)
Figure S2: Logarithm of the radial probability density $\ln|\psi(r)|^{2}$ of a representative Anderson-localized state as a function of the radial distance $r$ from the localization center, for a system with $L=5$ and $\langle A\rangle/V_{0}=0.3$. The linear behavior on the logarithmic scale confirms the exponential spatial decay characteristic of Anderson localization.
Figure S3: (Left) $\text{IPR}_{2}$ versus energy comparing clean and disordered systems, and (right) $\text{IPR}_{2}$ versus energy for different system sizes, showing progressively narrower absolute energy interval as the system size increases .
=======\!=\!=\!=\!=\!=\; [II] Deriving the Tight-Binding Anderson Model from a Continuum Potential
We start from the continuum single-particle Hamiltonian confined in a finite region:
$H=-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\mathrm{ext}}(\mathbf{r})+V_{\mathrm{imp}}(\mathbf{r}),$ (S-3)
where $V_{\mathrm{ext}}(\mathbf{r})$ is a periodic array of deep wells and $V_{\mathrm{imp}}(\mathbf{r})$ represents disorder modeled by Gaussian bumps:
$V_{\mathrm{imp}}(\mathbf{r})=\sum_{i}A_{i}\exp\!\left[-\frac{|\mathbf{r}-\mathbf{r}_{i}|^{2}}{2\sigma^{2}}\right].$ (S-4)
Assume each well in $V_{\mathrm{ext}}$ is deep and narrow, so the lowest eigenstate in each well is strongly localized. Let $\{\phi_{i}(\mathbf{r})\}$ denote these localized orbitals centered at lattice sites $\mathbf{R}_{i}$. In the deep-well limit:
$\phi_{i}(\mathbf{r})\approx\varphi(\mathbf{r}-\mathbf{R}_{i}),\quad\text{with negligible overlap between distinct wells.}$
The full wavefunction is expanded as:
$\Psi(\mathbf{r})=\sum_{i}c_{i}\phi_{i}(\mathbf{r}).$ (S-5)
The matrix elements in this basis are:
$\langle\phi_{i}|H|\phi_{j}\rangle$ $=\int d^{2}r\,\phi_{i}^{*}(\mathbf{r})\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\mathrm{ext}}(\mathbf{r})+V_{\mathrm{imp}}(\mathbf{r})\right]\phi_{j}(\mathbf{r}).$ (S-6)
For $i=j$ (on-site term):
$\epsilon_{i}=\langle\phi_{i}|H|\phi_{i}\rangle\approx E_{0}+\int d^{2}r\,|\phi_{i}(\mathbf{r})|^{2}V_{\mathrm{imp}}(\mathbf{r}),$ (S-7)
where $E_{0}$ is the ground-state energy of an isolated well. Since $V_{\mathrm{imp}}$ consists of Gaussian bumps and we assume one bump per well with negligible overlap, the integral picks up only the local impurity amplitude:
$\epsilon_{i}\approx E_{0}+A_{i}.$
For $i\neq j$ (hopping term):
$t_{ij}=\langle\phi_{i}|H|\phi_{j}\rangle\approx-\int d^{2}r\,\phi_{i}^{*}(\mathbf{r})\frac{\hbar^{2}}{2m}\nabla^{2}\phi_{j}(\mathbf{r}),$ (S-8)
which is exponentially small due to negligible overlap. We retain only nearest-neighbor hoppings $t$.
Collecting terms, the effective Hamiltonian in the localized basis is:
$H_{\mathrm{TB}}=\sum_{i}\epsilon_{i}c_{i}^{\dagger}c_{i}+\sum_{\langle i,j\rangle}t\left(c_{i}^{\dagger}c_{j}+c_{j}^{\dagger}c_{i}\right),$ (S-9)
with
$\epsilon_{i}=E_{0}+A_{i},\quad t\approx\text{constant (nearest-neighbor)}.$
Since $A_{i}$ are random and uncorrelated in the deep-well limit, $\epsilon_{i}$ are random on-site energies. Thus, Eq. (S-9) is the standard 2D Anderson Hamiltonian:
$H_{\mathrm{Anderson}}=\sum_{i}\epsilon_{i}c_{i}^{\dagger}c_{i}+t\sum_{\langle i,j\rangle}\left(c_{i}^{\dagger}c_{j}+\text{h.c.}\right),$ (S-10)
where $\epsilon_{i}$ are drawn from a disorder distribution set by the Gaussian amplitudes $A_{i}$.
Key Assumptions:
(1) Deep wells $\Rightarrow$ localized orbitals with negligible overlap.
(2) One Gaussian bump per well $\Rightarrow$ uncorrelated on-site disorder.
(3) Retain only nearest-neighbor hopping $\Rightarrow$ standard TB connectivity.
This derivation shows that the continuum model with isolated Gaussian impurities reduces to the tight-binding Anderson model in the deep-well limit.